how can i graph a linear system and tell how many solutions the system has
9x-3y=18
-12x+ 4y=24

Answer by richard1234(7193)

Multiplying the first equation by 4 and the second by 3, we get

36x - 12y = 72
-36x + 12y = 72

Adding them, we get 0 = 144. This means that if there is a solution (x,y) then 0 = 144. Therefore, no solutions.